OpenStudy (anonymous):

Using DeMoivre's Theorem, what is the value of [3(cos60* + i sin 60*)]?

OpenStudy (anonymous):

this is not demoivre, this is a direct calculation

OpenStudy (anonymous):

\[\cos(60)=\frac{1}{2}, \sin(60)=\frac{\sqrt{3}}{2}\] and so you have \[\frac{3}{2}+\frac{3\sqrt{3}}{2}i\]

OpenStudy (anonymous):

if you wanted to for example raise this number to the power of 3, then you would use demoivre